OpenStudy (anonymous):

During the first part of a trip, a canoeist travels 68 miles at a certain speed. The canoeist travels 14 miles on the second part of the trip at a speed 5 mph slower. The total time for the trip is 3 hours. What was the speed on each part of the trip?

OpenStudy (anonymous):

Vavg=68+14/3=27.33mi/h v2=v1-5 V2+V1/2=Vavg 2V1-5=54.66 V1=29.833mi/h V2=24.833mi/h